OpenStudy (anonymous):

A mountain has an elevation of 19,293 feet. In 1901, the glacier on this peak covered 4 acres. By 2005 this glacier had melted to only 1 acre. A) Assume that this glacier melted at a constant rate each year. Find this yearly rate. B) Use your answer from (A) to write a linear equation that gives the acreage A of this glacier t years past 1901.

OpenStudy (allank):

The melting of the glacier can fit really nicely onto an exponential decay function of the form: q=Q*e^(-kt) where q=amount remaining, Q=initial amount, e=exponential, k=melting rate, t=time in years. Note that we use -k because the amount of ice is reducing. So from the problem, Q=4 acres, q=1 acre, e=value in calculator of course, t=(2005-1901)=104 years. Plug those into the equation, 1=4*e^(-104k) Now solve for k to get the constant rate of melting. I hope that helps.

OpenStudy (anonymous):

so what about B?

OpenStudy (anonymous):

i'm still kind of lost

OpenStudy (allank):

Alright. Have you learnt about exponential functions?

OpenStudy (anonymous):

A little

OpenStudy (allank):

I have applied the that knowledge here. Exponential functions have the form: Q=Q(1)*e^(-kt) Where Q is the quantity being studied, and Q(1) is the initial amount of the quantity, e is the exponential(value given in calculator), and k is the constant of decay or growth (in this case decay because it is negative, and t is the time elapsed. Does this sound familiar?

OpenStudy (anonymous):

no not at all! uggghhhh

OpenStudy (allank):

Alright. We can still do this. Which math class are you taking, so that I know which methodology to explain with.

OpenStudy (anonymous):

university of phoenix mat/116

OpenStudy (allank):

Alright. Which category does that fall into? Calculus or Algebra?

OpenStudy (anonymous):

algebra

OpenStudy (allank):

No prob. You'll get this really fast. Just a sec...

OpenStudy (allank):

Now, imagine that we have a certain car whose value decreases by 5% each month. If its initial value 2 months ago was $1000, 1 month ago its value was 95% of 1000=950, last month its value was 95% of 950 = 902.5, this month its value is 95% of 902.5=857.375. Notice the trend: 1000,950,902.5,857.375. This trend fits into the equation: value=initial value*percentage remaining^time e.g. value now = 1000*.95^3 =857.375 From this case scenario, we can derive the formula: amount=initial amount*percentage remaining^time. Can you use this formula to solve the question now?

OpenStudy (anonymous):

i'll try if not I'll just skip it

OpenStudy (anonymous):

thanks

OpenStudy (allank):

You're welcome.